# RouterArena: An Open Platform for Comprehensive Comparison of LLM Routers

### v1

@@ Abstract
Today’s LLM ecosystem comprises a wide spectrum of models that differ in size, capability, and cost. No single model is optimal for all scenarios; hence, LLM routers have become essential for selecting the most appropriate model under varying circumstances. However, the rapid emergence of various routers makes choosing the right one increasingly challenging. To address this problem, we need comprehensive router comparison and a standardized leaderboard, similar to those available for models. In this work, we introduce RouterArena, the first open platform enabling comprehensive comparison of LLM routers. RouterArena has (1) a principally constructed dataset with broad knowledge domain coverage, (2) distinguishable difficulty levels for each domain, (3) an extensive list of evaluation metrics, and (4) an automated framework for leaderboard updates. Leveraging our framework, we have produced the initial leaderboard with detailed metrics comparison as shown in Figure 1. We will soon open our platform to the public.
@@ 1 Introduction
Given its importance and promise, many LLM routers have recently emerged in both industry and academia (Figure 2). A notable example is GPT-5 [27], which incorporates routing as a key feature by directing user queries to the most suitable model within the OpenAI family. As routers proliferate, the challenge shifts from selecting the right model to selecting the right router. Unfortunately, router evaluation has not kept pace: there is currently no open evaluation platform, akin to LMArena [5], that systematically compares open-source routers [15, 42] and commercial routing services [25, 24] under a unified protocol.
@@ 2 Motivation
As shown in Figure 2, the landscape of LLM routers is rapidly expanding, evolving from academic exploration to commercial deployment. From a few scattered academia routers [39, 2, 13] in mid-2023, the number of publications expanded to more than a dozen by 2024 [22, 43, 3, 41]. By 2025, not only did academia routers continue to grow [36, 16, 9, 40], but commercial products also emerged [25, 24], most notably GPT-5 [27] with a built-in router.
@@
Table 1: Comparison of existing works [15, 17, 11, 42] and RouterArena. RouterArena enables comprehensive router comparison with extensive query categories, difficulty levels, evaluation metrics, and router inclusion.
@@ 3 RouterArena Evaluation Dataset
To evaluate a router’s ability to recognize problem domains and route queries to the appropriate specialist models, the dataset must provide broad domain coverage. To achieve this, we draw inspiration from the Dewey Decimal Classification (DDC) system [8], a book classification framework widely used in libraries. The DDC is renowned for its comprehensive and logical structure, providing a proven methodology for organizing the entire world of knowledge into distinct, hierarchical categories [30].
Principle 2: Bloom-Based Distinguishable Difficulty Levels.
To evaluate whether a router can determine query difficulty and make accuracy-cost tradeoffs—choosing between powerful but expensive models and weaker but cheaper ones—the dataset must include clearly distinguishable difficulty levels. To structure these difficulty levels, we adopt Bloom’s taxonomy [1], a widely used framework in quantifying question complexity [34, 14, 28]. It works by classifying cognitive tasks into six ascending categories: remembering, understanding, applying, analyzing, evaluating, and creating. In this work, we further group these six levels into easy-medium-difficult three levels.
@@
Following these two principles, we curate our evaluation dataset as follows. To ensure coverage across all DDC categories (excluding religion), we first collect all the queries from two existing LLM benchmark datasets and then supplement underrepresented categories with data from 21 open-source, domain-specific datasets. To determine the difficulty level of each query based on Bloom’s taxonomy, we employ an LLM-as-Judge approach with DeepSeek-V3.1 [7] (prompt specified in Appendix D), enabling automatic difficulty annotation. We exclude the create-type questions because they are open-ended and cannot be reliably evaluated. We then categorize remembering and understanding questions as easy, applying questions as medium, and analyzing and evaluating questions as difficult.
Next, to fairly distribute questions across categories and difficulty levels, we propose a recursive deficit redistribution algorithm. We begin by setting the ratio of science to humanities at 2:1. Within each top-level category, if a sub-category falls short of its proportional quota, the resulting surplus is recursively and uniformly redistributed to those sub-categories that exceed their initial allocation. We apply the same procedure within each sub-category to allocate data across different difficulty levels, ensuring balanced coverage throughout the dataset.
The above steps yield approximately 62,000 queries in total. However, this raw dataset contains many highly similar or even duplicate questions inherited from the various sampled sources. Such redundancy does not benefit router evaluation and may even introduce noise into the results. To address this, we perform cosine-similarity–based de-duplication using sentence-transformers/all-MiniLM-L6-v2. By strictly following the allocation strategy and selecting the least similar samples, we ensure that the resulting RouterArena dataset maintains broad coverage with minimal redundancy.
@@
Our final evaluation dataset consists of 8,400 queries sampled from 23 source datasets. It spans nine top-level domains and 44 categories, with each category containing queries across three difficulty levels. Figure 3 illustrates the detailed composition of the dataset. Note that the distribution of difficulty levels is skewed, but it reflects real-world query patterns—easy and medium questions occurring more frequently than hard ones. We include more details about the dataset in Appendix B.1, including dataset schema and concrete query examples.
@@ 5.1 Arena Ranking
RouterArena provides a series of router leaderboards that enables users to compare the capabilities of different routers and select the one best suited to their scenarios. It includes six ranking scores based on the evaluation metrics described in Section 4, including Arena, Optimal-selection-ratio, Optimal-acc-ratio, Optimal-cost-ratio, Robustness, and Latency. Among these, the Arena score captures the trade-off between accuracy and cost by combining them into a single composite measure using the Weighted Harmonic Mean [12]. Specifically, to better distinguish between routers with low costs, we apply a base-2 logarithmic ($log_{2}$) transformation to the cost values. Under this scaling, each doubling of price reduces the cost score by one unit. For router $i$ with cost $c_{i}$, we define its normalized cost as
@@
where the parameter $\beta>0$ controls the relative importance of accuracy versus cost. Setting $\beta>1$ places greater weight on cost. By default, we use $\beta=0.1$, emphasizing routing accuracy, because highly accurate routers are generally more valuable even if they incur slightly higher costs.
@@ 6.1 Experimental Settings
For commercial routers, we evaluated the router from Not Diamond [25], which provides access to over 60 models, and the Azure Model Router [24], which currently only supports OpenAI models. We also included GPT-5 [27], whose model family incorporates an internal router. For NotDiamond, we selected 26 representative models spanning different parameter scales, architectures, and reasoning abilities. For Azure-Router, we evaluated the entire model pool, including GPT-5 model families. Appendix A provides full details of the model pools used for each router.
For open-source routers, we evaluated nine representative systems covering a diverse routing approaches. Specifically, we chose both the KNN- and MLP-based methods trained on RouterBench [15] as baselines. We further included GraphRouter [10], which leverages graph neural networks (GNNs) for routing, and the Universal Router [19], which uses K-means clustering. To capture cost–accuracy tradeoffs, we evaluated CARROT Router [31], while RouterDC [4] was incorporated as a dual contrastive learning–based approach. Additionally, we considered IRT-Router [32], which applies item response theory to explicitly model the interaction between query attributes and model capabilities, and RouteLLM [26], which performs binary selection between a stronger and a weaker model. Moreover, we also take the latest vLLM Semantic Router [35] into consideration, which leverages a ModernBERT [18] to categorize the incoming requests into pre-defined categories, and selects the model that has the highest score.
@@ 6.2 Results
Figure 6: Normalized Deferral Curve
@@
Given that user prompts are often noisy, we further assess router sensitivity and robustness. Specifically, we prepend an irrelevant keyword to the input and observe whether the router alters its original model selection. We define robustness as 1 – the proportion of changed selections. As shown in Figure 8, routers leveraging latent representations of prompts demonstrate stronger stability against noise, whereas methods relying on explicit representations, such as KNN and NIRT-BERT, are considerably more sensitive. These findings highlight the importance of applying prompt engineering techniques to mitigate noisy queries in the future.
@@ 7 Related Work
With the increasing availability of specialized models that can surpass even the most capable general-purpose LLMs in specific domains, both academia and industry have been actively exploring how to build LLM routers. In industry, several systems have emerged. Martian Router [37] proposed the idea of model mapping, while Storytell [33] categorizes user queries and routes them to the best-performing models. Other companies also seek to find the optimal model for user’s tasks by balancing performance and cost [25, 29, 27]. Recent academic efforts have also begun to emerge. GraphRouter [10] leverages graph neural networks, and Router-R1 [38] employs reinforcement learning. The growth of open-source solutions underscores the need for effective router evaluation [31, 32, 11].
LLM Router Benchmark.
RouterBench [15] introduces a large-scale dataset consisting of over 405k inference outcomes from representative LLMs. RouterEval [17] collects performance results from 8,500 LLMs across 12 widely used benchmarks. FusionBench [11] covers 14 tasks across five domains and leverages 20 open-source LLMs. Other benchmarks have also contributed to this line of work by using different data collection methods [20, 23]. However, these benchmarks fail to provide broad coverage across disciplinary domains or cover all kinds of routers. In contrast, our benchmark is systematically constructed based on an authoritative knowledge classification framework, making it the first comprehensive and actionable benchmark for LLM routing.
@@ Appendix A Model Pools by Router
We provide the model pool used by each router in Table 3. Each router is trained on its own distinct set of LLMs. When replicating training runs, the same LLM allocation is preserved for each router to maintain consistency in usage. Since some models do not have a serverless inference endpoint, we utilized vLLM [21] to perform inference on them using an 8x A100 40 GB server. The Together.ai price bracket in Fig 4 estimates the inference price based on the parameter size.
Dense Models Mixture-of-Experts Models
Model Size Price Model Size Price
Up to 4B \$0.10 Up to 56B \$0.60
4.1B – 8B \$0.20 56.1B – 176B \$1.20
8.1B – 21B \$0.30 176.1B – 480B \$2.40
21.1B – 41B \$0.80
41.1B – 80B \$0.90
80.1B – 110B \$1.80
Appendix B RouterArena Dataset
B.1 Datasets Details
Figure 9: The composition datasets in RouterArena.
@@ B.1 Datasets Details
Note, we opt out of the domain of Religion due to its sensitivity in its content of the topic.
@@
We provide detailed dataset columns in Table 5. Domain and Category represent the higher and lower classification levels defined by Dewey Decimal Classification (DDC) [8]. Dataset Name represents the specific dataset that the question comes from. Global Index is a unique identifier for each question across the benchmark dataset. Context, Question, and Options are question-specific fields, and may be used to provide more information to curate the input question after formatting it into the input prompt. Answer is the golden answer. Bloom Level is the classified difficulty level based on Bloom’s Taxonomy [1].
Figure 10: The difficulty distribution in RouterArena.
Figure 10 shows the overall difficulty distribution in our benchmark dataset.
B.2 Dataset Examples
@@
Figure 11 shows some examples from the benchmark dataset.
Appendix C Leaderboard

### v3

@@ Abstract
Today’s LLM ecosystem comprises a wide spectrum of models that differ in size, capability, and cost. No single model is optimal for all scenarios; hence, LLM routers have become essential for selecting the most appropriate model under varying circumstances. However, the rapid emergence of various routers makes choosing the right one increasingly challenging. To address this problem, we need comprehensive router comparison and a standardized leaderboard, similar to those available for models. In this work, we introduce RouterArena, the first open platform enabling comprehensive comparison of LLM routers. RouterArena has (1) a principally constructed dataset with broad knowledge domain coverage, (2) distinguishable difficulty levels for each domain, (3) an extensive list of evaluation metrics, and (4) an automated framework for leaderboard updates. Leveraging our framework, we have produced the initial leaderboard with detailed metrics comparison as shown in Figure 1. Our framework for evaluating new routers is on https://github.com/RouteWorks/RouterArena. Our leaderboard is on https://routeworks.github.io/
@@ 1 Introduction
Given its importance and promise, many LLM routers have recently emerged in both industry and academia (Figure 2). A notable example is GPT-5 [openai2025gpt5], which incorporates routing as a key feature by directing user queries to the most suitable model within the OpenAI family. As routers proliferate, the challenge shifts from selecting the right model to selecting the right router. Unfortunately, router evaluation has not kept pace: there is currently no open evaluation platform, akin to LMArena [chiang2024chatbot], that systematically compares open-source routers [routerbench, embedllm] and commercial routing services [notdiamond, azure-model-router-2025] under a unified protocol.
@@ 2 Motivation
As shown in Figure 2, the landscape of LLM routers is rapidly expanding, evolving from academic exploration to commercial deployment. From a few scattered academia routers [zhang2023model, chen2023frugalgpt, hari2023tryage] in mid-2023, the number of publications expanded to more than a dozen by 2024 [liu2024optllm, zhuang2024embedllm, chen2024routerdc, zhao2024texttt]. By 2025, not only did academia routers continue to grow [wang2025mixllm, huang2025routereval, ding2025best, zhang2025capability], but commercial products also emerged [notdiamond, azure-model-router-2025], most notably GPT-5 [openai2025gpt5] with a built-in router.
@@
Table 1: Comparison of existing works [routerbench, routereval, fusion, embedllm] and RouterArena. RouterArena enables comprehensive router comparison with extensive query categories, difficulty levels, evaluation metrics, and router inclusion.
@@ 3 RouterArena Evaluation Dataset
To evaluate a router’s ability to recognize problem domains and route queries to the appropriate specialist models, the dataset must provide broad domain coverage. To achieve this, we draw inspiration from the Dewey Decimal Classification (DDC) system [DDC], a book classification framework widely used in libraries. The DDC is renowned for its comprehensive and logical structure, providing a proven methodology for organizing the entire world of knowledge into distinct, hierarchical categories [satija2013theory].
Principle 2: Bloom-Guided Cognitive Skill Levels.
Routers in practice face queries that range from simple factual recall to complex reasoning. To ensure coverage of cognitive skills from simple recall to higher-order reasoning and judgment, we adopt Bloom’s taxonomy [bloom1956taxonomy], a widely used framework for quantifying question complexity and cognitive processes [ddc1, ddc2, ddc3].
Principle 3: Empirically Defined Difficulty Levels.
To test whether a router can make meaningful accuracy–cost tradeoffs—choosing between powerful but expensive models and weaker but cheaper ones—the dataset must also provide clearly distinguishable difficulty levels. We therefore define difficulty empirically: for each query, we compute the number of models (out of 42) that answer it correctly and sort queries by this empirical difficulty.
@@
Following these two principles, we curate our evaluation dataset as follows. To ensure coverage across all DDC categories (excluding religion), we first collect all the queries from two existing LLM benchmark datasets and then supplement underrepresented categories with data from 21 open-source, domain-specific datasets, resulting in 62,000 raw queries. We then cap each source dataset at approximately 1,000 queries. [^1] It primarily affects large corpora such as WMT and SuperGLUE and yields roughly 30,000 candidate queries while preventing a few sources from dominating the pool To determine the cognitive skill level of each query based on Bloom’s taxonomy, we employ an LLM-as-Judge approach with DeepSeek-V3.1 [DS-V3.1] (prompt specified in Appendix G). We use these Bloom levels only to characterize the cognitive skills involved in each query, not as ground-truth difficulty labels, and we further validate them with a small-scale human study (Appendix B.1).
Next, to fairly distribute questions across categories and cognitive levels, we propose a recursive deficit redistribution algorithm. We begin by setting the ratio of science to humanities at 2:1. Within each top-level category, if a sub-category falls short of its proportional quota, the resulting surplus is recursively and uniformly redistributed to those sub-categories that exceed their initial allocation. We apply the same procedure within each sub-category to allocate data across different cognitive levels, ensuring balanced coverage throughout the dataset.
However, this raw dataset contains many highly similar or even duplicate questions inherited from the various sampled sources. Such redundancy does not benefit router evaluation and may even introduce noise into the results. To address this, we perform cosine-similarity–based de-duplication using sentence-transformers/all-MiniLM-L6-v2. By strictly following the allocation strategy and selecting the least similar samples, we ensure that the resulting RouterArena dataset maintains broad coverage with minimal redundancy.
@@
Our final evaluation dataset consists of 8,400 queries sampled from 23 source datasets. It spans nine top-level domains and 44 categories in Figure 3 (Left). We define difficulty empirically: for each query, we count how many of the 42 models answer it correctly and group queries into three bands—hard ($\leq$4/42), medium (5–19/42), and easy ($\geq$20/42)—as shown in Figure 3 (right). We provide the model list of these 42 models and the rationale to use them at Appendix B.4. Figure 4 further shows that, when sorted by this measure, both the overall and per-domain curves increase smoothly from low to high accuracy without large gaps, indicating a well-spread spectrum of difficulty that can distinguish router behavior. Note that this empirical difficulty distribution is skewed, but it reflects real-world query patterns—easy and medium questions occurring more frequently than hard ones. We include more details about the dataset in Appendix B.2, including dataset schema and concrete query examples.
Figure 4: Each query is answered by 42 LLMs and evaluated for accuracy. Sorting queries by empirical accuracy yields smooth, monotonic curves overall and within each domain, indicating that query difficulty is well spread across the full range from very hard (left) to very easy (right).
@@ 5.1 Arena Ranking
RouterArena provides a series of router leaderboards that enables users to compare the capabilities of different routers and select the one best suited to their scenarios. It includes six ranking scores based on the evaluation metrics described in Section 4, including Arena Score, Optimal-selection-ratio, Optimal-acc-ratio, Optimal-cost-ratio, Robustness, and Latency. Among these, the Arena score $S_{i,\beta}$ captures the trade-off between accuracy and cost by combining them into a single composite measure using the Weighted Harmonic Mean [w_h_m]. Specifically, to better distinguish between routers with low costs, we apply a base-2 logarithmic ($log_{2}$) transformation to the cost values. Under this scaling, each doubling of price reduces the cost score by one unit. For router $i$ with cost $c_{i}$, we define its normalized cost as
@@
where the parameter $\beta>0$ controls the relative importance of accuracy versus cost. Intuitively, $\beta$ specifies the relative weighting applied to cost: we assign cost a weight of $\beta$ in comparison to accuracy. By default, we use $\beta=0.1$, emphasizing routing accuracy, because highly accurate routers are generally more valuable even if they incur slightly higher costs.
Details about how we update $c_{min}$ and $c_{max}$, recommendations on values of $\beta$ for different user needs are provided in Appendix F.
@@ 6.1 Experimental Settings
For commercial routers, we evaluated the router from Not Diamond [notdiamond], which provides access to over 60 models, and the Azure Model Router [azure-model-router-2025], which currently only supports OpenAI models. We also included GPT-5 [openai2025gpt5], whose model family incorporates an internal router. For NotDiamond, we selected 26 representative models spanning different parameter scales, architectures, and reasoning abilities. For Azure-Router, we evaluated the entire model pool, including GPT-5 model families. Appendix A provides full details of the model pools used for each router.
For open-source routers, we evaluated nine representative systems covering a diverse routing approaches. Specifically, we chose both the KNN- and MLP-based methods trained on RouterBench [routerbench] as baselines. We further included GraphRouter [graphrouter], which leverages graph neural networks (GNNs) for routing, and the Universal Router [universal], which uses K-means clustering. To capture cost–accuracy tradeoffs, we evaluated CARROT Router [carrot], while RouterDC [routerdc] was incorporated as a dual contrastive learning–based approach. Additionally, we considered IRT-Router [irtrouter], which applies item response theory to explicitly model the interaction between query attributes and model capabilities, and RouteLLM [routellm], which performs binary selection between a stronger and a weaker model. Moreover, we also take the latest vLLM Semantic Router [vllm_semantic_router] into consideration, which leverages a ModernBERT [modernbert] to categorize the incoming requests into pre-defined categories, and selects the model that has the highest score.
@@ 6.2 Results
Figure 7: Normalized Scatter Plot
@@
Given that user prompts are often noisy, we further assess router sensitivity and robustness. The detailed evaluation is illustrated in Appendix E. We define robustness as 1 – the proportion of changed selections. As shown in Figure 9, robustness is uniformly low across both commercial routers and academic routers such as vLLM-SR, MIRT-BERT, and NIRT-BERT. A key reason is that these academic routers depend on BERT to compute latent representations, and BERT is known to be sensitive to surface-level perturbations [bert_robust]. These findings highlight the importance of applying prompt engineering techniques to mitigate noisy queries.
@@
Generation Length
We also report the average generation length and price per M output tokens for each router in Table 8. GPT-5 and Azure-Router rely more heavily on reasoning models, leading to substantially longer generations than most other routers. Longer answers are not inherently better; this metric simply offers an additional axis for users to trade off concise versus detailed responses.
6.3 Results by Difficulty Levels
Table 6 shows the accuracy–cost tradeoff across easy, medium, and hard queries. Most routers exceed 89% accuracy on easy questions, but accuracy drops sharply on medium and especially hard ones (often below 10%), confirming that our difficulty bands are non-trivial. GPT-5 achieves the highest accuracy at every level, but at a much higher cost. At the same time, routers such as Azure-Router and vLLM-SR offer competitive accuracy on easy/medium queries at a lower cost. This contrast reveals substantial headroom on hard questions and clear efficiency–accuracy tradeoffs across methods. We also observe that some routers (e.g., GPT-5, Azure-Router, MIRT-BERT, NotDiamond) allocate significantly more budget to Hard than Easy queries, while others show almost flat cost across difficulty, suggesting room for more difficulty-aware routing strategies.
6.4 Additional Results on LongBench-v2
We provide additional results on LongBench-v2 in Appendix C. On this long-context suite, GPT-5 attains the highest accuracy, while Azure-Router and NIRT-BERT offer competitive performance at substantially lower cost, illustrating clear accuracy–cost trade-offs in the long-context regime.
@@ 7 Related Work
With the increasing availability of specialized models that can surpass even the most capable general-purpose LLMs in specific domains, both academia and industry have been actively exploring how to build LLM routers. In industry, several systems have emerged. Martian Router [martian] proposed the idea of model mapping, while Storytell [storytell] categorizes user queries and routes them to the best-performing models. Other companies also seek to find the optimal model for user’s tasks by balancing performance and cost [notdiamond, requestyai, openai2025gpt5]. Recent academic efforts have also begun to emerge. GraphRouter [graphrouter] leverages graph neural networks, and Router-R1 [routerr1] employs reinforcement learning. The growth of open-source solutions underscores the need for effective router evaluation [carrot, irtrouter, fusion].
LLM Router Benchmarks.
RouterBench [routerbench] introduces a large-scale dataset consisting of over 405k inference outcomes from representative LLMs. RouterEval [routereval] collects performance results from 8,500 LLMs across 12 widely used benchmarks. FusionBench [fusion] covers 14 tasks across five domains and leverages 20 open-source LLMs. Other benchmarks have also contributed to this line of work by using different data collection methods [dsc, omnirouter]. All of these benchmarks adopt an offline design: they release a large frozen corpus of model outputs that is ideal for cheap, fully reproducible comparison of routing algorithms under a fixed model pool. By contrast, RouterArena is explicitly designed as a live leaderboard that includes both academic and commercial routers. Our live leaderboard, by nature, encourages router designers to explore evolving and specialized models, making routing decisions more meaningful in practice. Meanwhile, RouterArena improves on prior router benchmarks by covering more tasks and domains, and by providing a dataset with empirically calibrated difficulty, together with more comprehensive metrics.
@@ Appendix A Model Pools by Router
We provide the model pool used by each router in Table 3. Each router is trained on its own distinct set of LLMs. When replicating training runs, the same LLM allocation is preserved for each router to maintain consistency in usage. Since some models do not have a serverless inference endpoint, we utilized vLLM [vllm_kwon2023efficient] to perform inference on them using an 8x A100 40 GB server. The Together.ai price bracket in Fig 9 estimates the inference price based on the parameter size.
Appendix B Datasets Details
B.1 Use of LLM-as-a-judge
We sampled 450 queries from the dataset by stratifying along both the domain and Bloom level, ensuring that each combination was proportionally represented. Eighteen volunteers (undergraduate and graduate students) were asked to assign a Bloom level to each query under the same instructions as used for the DeepSeek-V3.1 judge. For each query, we compare the human label with the LLM’s prediction. We observe 54.9% exact agreement and 76.7% agreement within $\pm$1 Bloom level, indicating that most disagreements occur between adjacent levels in this inherently subjective taxonomy. Since Bloom labels are used only to ensure cognitive-skill coverage during dataset construction—while all evaluation difficulty is defined empirically via accuracy over 42 LLMs—this level of agreement is sufficient for our purposes.
B.2 Dataset composition and Schema
(a) Dataset source.
(b) Difficulty distribution.
Figure 10: Overview of dataset composition: (a) data sources and (b) difficulty distribution.
@@ B.2 Dataset composition and Schema
Table 4 shows the detailed dataset columns.
B.3 Dataset Examples
@@
Figure 11 shows some dataset examples.
Table 5: These are the 42 models used for empirical difficulty labeling. Models span across a range of sizes and performances, showcasing that RouterArena could distinguish LLMs by providing diverse questions of difficulty.
Rank Model Score [%]
1 gpt-5 $74.04$
2 claude-3-5-sonnet $67.73$
3 deepseek/deepseek-chat $66.86$
4 gpt-4o $66.35$
5 llama-3-1-405b-instruct $65.43$
6 glm-4-plus $62.02$
7 gpt-4-1106-preview $61.91$
8 meta-llama_Meta-Llama-3.1-70B-Instruct-Turbo $60.64$
9 Qwen/Qwen2.5-32B-Instruct-GPTQ-Int4 $60.51$
10 Qwen/Qwen2.5-72B-Instruct $59.98$
11 meta-llama_Llama-3.1-70B-Instruct $59.89$
12 Qwen/QwQ-32B $59.81$
13 gpt-4o-mini $59.22$
14 meta-llama_Meta-Llama-3-70B-Instruct $58.16$
15 google/gemini-flash-1.5 $57.21$
16 glm-4-air $54.63$
17 Qwen/Qwen1.5-72B-Chat $49.86$
18 Qwen/Qwen2.5-7B-Instruct $47.70$
19 glm-4-flash $47.62$
20 mistralai/Mixtral-8x7B-Instruct-v0.1 $47.01$
21 01-ai/Yi-34B-Chat $45.54$
22 llama-3-1-8b-instruct $45.23$
23 NousResearch/Nous-Hermes-2-Yi-34B $45.01$
24 meta-llama_Llama-3.1-8B-Instruct $44.24$
25 meta-llama_Llama-2-70b-chat-hf $39.82$
26 meta/llama-2-70b-chat $39.49$
27 mistralai/Ministral-8B-Instruct-2410 $32.87$
28 llama-3-2-3b-instruct $32.55$
29 cognitivecomputations/dolphin-2.9-llama3-8b $32.35$
30 Qwen/Qwen2.5-32B $32.26$
31 mistralai/Mistral-7B-Instruct-v0.2 $31.83$
32 meta-llama_llama-3-8b-instruct $30.45$
33 cognitivecomputations/dolphin-2.6-mistral-7b $29.98$
34 mistralai/Mistral-7B-Instruct-v0.3 $28.05$
35 meta-llama_Meta-Llama-3-8B-Instruct $23.69$
36 WizardLM/WizardLM-13B-V1.2 $23.03$
37 llama-3-2-1b-instruct $22.10$
38 meta-llama_Llama-2-7b-chat-hf $20.50$
39 meta/codellama-34b-instruct $20.42$
40 codellama/CodeLlama-34b-Instruct-hf $19.31$
41 meta-llama_Meta-Llama-2-7B $18.92$
42 meta-llama_Meta-Llama-2-7B-hf $18.39$
B.4 Model List for Empirical Difficulty Labeling
We use the following 42 models to compute the empirical difficulty labels shown in Figure 5. These models are drawn from the model pools of the routers we benchmark, and were chosen to cover a broad range of parameter scales, architectures, and providers. This diversity ensures that queries elicit a wide spread of accuracies across models, which we indeed observe in their scores and use as the basis for our empirical difficulty measure.
Appendix C Additional Results and Tables:
Table 6: Router performance by difficulty level (accuracy and cost per 1k queries).
Router Overall Easy Medium Hard
GPT-5 74.0% (\$14.02) 95.1% (\$5.68) 68.6% (\$14.80) 27.5% (\$35.73)
Azure-Router 68.1% (\$0.54) 93.3% (\$0.30) 59.5% (\$0.63) 17.9% (\$1.05)
NotDiamond 68.0% (\$9.34) 92.5% (\$6.34) 61.9% (\$10.29) 15.1% (\$15.16)
vLLM-SR 67.3% (\$1.67) 95.3% (\$1.56) 57.9% (\$1.70) 8.7% (\$1.87)
CARROT 67.2% (\$2.06) 95.0% (\$1.79) 58.0% (\$2.26) 9.0% (\$2.36)
MIRT-BERT 66.9% (\$0.15) 95.9% (\$0.10) 56.8% (\$0.17) 7.1% (\$0.26)
NIRT-BERT 62.0% (\$0.69) 93.7% (\$0.53) 47.0% (\$0.78) 4.7% (\$0.90)
MLP 61.6% (\$4.83) 93.2% (\$3.54) 46.6% (\$5.74) 4.6% (\$6.52)
KNN 58.7% (\$4.27) 89.2% (\$3.08) 43.6% (\$5.13) 4.5% (\$5.77)
Graphrouter 57.0% (\$0.34) 89.7% (\$0.25) 39.3% (\$0.39) 2.4% (\$0.46)
RouteLLM 47.0% (\$0.27) 79.8% (\$0.20) 24.4% (\$0.33) 2.0% (\$0.32)
RouterDC 32.0% (\$0.07) 54.9% (\$0.06) 15.3% (\$0.08) 2.2% (\$0.09)
Long-context routing.
Table 7 reports router performance on the LongBench-v2 dataset. We randomly sampled 100 queries of short and medium length from the dataset. GPT-5 attains the highest accuracy (71%) but also the highest cost per 1k queries. Azure-Router, MIRT-BERT, and NIRT-BERT achieve reasonably strong accuracy (60–67%) at one order of magnitude lower cost, whereas methods like MLP, KNN, NotDiamond, and vLLM-SR are substantially more expensive for comparable or worse accuracy. RouteLLM cannot be evaluated in this setting because its text encoder only supports inputs up to 8,192 tokens, which excludes many of our long-context queries.
Table 7: RouterArena routers—long-context performance and cost. RouteLLM cannot be evaluated on this subset because its text encoder only supports inputs up to 8,192 tokens.
Router Acc (%) Cost / 1K Queries (USD)
GPT-5 71 45.70
NotDiamond 70 59.30
Azure-Router 67 9.54
NIRT-BERT 60 6.60
RouteLLM
Graphrouter 42 23.60
vLLM-SR 42 64.80
CARROT 25 4.10
RouterDC 12 3.60
MIRT-BERT 60 7.50
MLP 65 286.70
KNN 69 300.30
Generation length and cost.
Table 8 summarizes average generation length and cost per million output tokens. GPT-5 produces the longest responses and is also the most expensive per token, while Azure-Router, NIRT-BERT, and MIRT-BERT generate shorter outputs at much lower cost. In contrast, heuristic baselines such as MLP, KNN, and CARROT have relatively short generations yet very high token-level costs, indicating that they are inefficient choices when generation cost is a primary concern.
Table 8: RouterArena routers—average generation length and cost.
Router Avg. Gen. Length Cost / M Tokens (USD)
GPT-5 970.65 10.00
NotDiamond 390.05 9.18
Azure-Router 270.90 1.79
NIRT-BERT 215.82 1.33
RouteLLM 140.24 0.68
Graphrouter 132.24 0.81
vLLM-SR 109.27 9.03
CARROT 105.56 12.58
RouterDC 102.83 0.20
MIRT-BERT 89.16 1.00
MLP 81.06 29.75
KNN 79.80 26.31
Appendix D Leaderboard